# Bayesian reconstruction of primordial perturbations from induced gravitational waves

### v2

@@ 3.2 Bayesian Model Selection via Bayesian Evidence
[…]
where $\mathcal{L}(d|\theta,\mathcal{M})$ is the likelihood of the data $d$ given the parameters $\theta$ and model $\mathcal{M}$, and $\pi(\theta|\mathcal{M})$ is the prior distribution of the model parameters [34]. Thus, for each spline model, the parameters $\theta$ over which we integrate are the spline node locations $\log_{10}k_{i}$, and the corresponding amplitudes $\log_{10}A_{i}$ of the scalar spectrum.
The Bayesian evidence automatically incorporates a penalty for excessive model complexity, effectively implementing Occam’s razor. In fact, a model with too many spline nodes has a larger parameter space to integrate over, diluting the integral and leading to a lower evidence unless the additional complexity is justified by the data. By comparing $\mathcal{Z}$ across different values of $N$, we select the models that best balance flexibility and predictive power. Typically, differences of $\Delta\log\mathcal{Z}\gtrsim 5$ between two models $(1,2)$ suggests that model $(1)$ is decisively favoured over model $(2)$, with $\Delta\log\mathcal{Z}_{(12)}=\log\mathcal{Z}_{(1)}-\log\mathcal{Z}_{(2)}$, while smaller differences ($\Delta\log\mathcal{Z}\lesssim 1$) are considered to be inconclusive in terms of Jeffrey’s scale [34].
The posterior distribution for the spline parameters can be highly complex, featuring multiple local maxima, as well as strong parameter degeneracies. Traditional sampling methods such as Markov Chain Monte Carlo (MCMC) struggle in such cases due to the difficulty of efficiently exploring the high-dimensional space. Additionally, local optimization methods are inadequate because they can become trapped in local extrema. For the same reasons, we opt not to use AIC or other frequentist criteria for model comparison. To overcome these challenges, we employ nested sampling [35], a particularly well-suited approach for computing Bayesian evidence while also generating posterior samples as a byproduct.
@@
[…]
Hence, $\bar{\mathcal{L}}$ can be thought of as the enclosing likelihood of cumulative prior mass $X$ which encompasses the region where the likelihood is larger than a threshold $\bar{\mathcal{L}}$. The core algorithm begins with a set of points drawn from the prior, called live points. At each step of the algorithm, the live point with the lowest likelihood is replaced by sampling from the prior, subject to the constraint that the likelihood at the new point must be larger than the current lowest likelihood value, and a statistical estimate of the prior mass $X$ is calculated. At the end of this process, we are left with a set of samples $\{\mathcal{L},X\}$ which can be used to estimate the evidence as well as approximate the posterior distribution of the parameters $\theta$ [36, 37]. This method has several advantages over MCMC: in addition to providing an estimate of the evidence, it can efficiently handle multi-modal posteriors, and a complex posterior geometry. We implement nested sampling using the publicly available libraries PolyChord [38, 39] and Nautilus [40].
@@ 3.3 Functional posteriors
[…]
@@ 4 Results
As an example, we choose the pivot frequency $f_{*}=10^{-3}\,\mathrm{Hz}$ as representative for LISA, with low uncertainties around the peak sensitivity $f_{*}$ and increasing ones towards the two ends of LISA’s sensitivity range (see e.g. [13]). We then use the result to create a Gaussian likelihood for the $\Omega_{\rm GW}$ as an input for the nested sampler, i.e.
@@
In our analysis, we assume that such $\Omega_{\mathrm{GW}}$ spectrum is reconstructed from detector data – e.g. using codes such as SGWBinner [41, 42] in the case of LISA – thus providing datasets consisting of the observed values of $\Omega_{\mathrm{GW}}$ and their associated uncertainties $\Delta\Omega_{\mathrm{GW}}$, over certain frequency ranges specific to the instruments. [^6] This assumption helps us to focus our attention entirely on our reconstruction method but is not a necessity. This reconstruction method (or any other) can be easily incorporated into existing codes in such a manner that eliminates the need for a separate initial inference step to obtain $\Omega_{\mathrm{GW}},\,\Delta\Omega_{\mathrm{GW}}$.
While the spectra and error bars are highly idealised, and artificially generated with LISA as representative experiment to keep in mind, the methods we present do not rely on this being the case. They are applicable for any general form for $\Omega_{\mathrm{GW}}$ and any GW experiment. In particular, one may consider more realistic realizations of the $\Omega_{\mathrm{GW}}$ data simulated using the LISA noise model, without changing our method  [^7] We should note though that the curvature spectrum reconstruction can only be as good as the $\Omega_{\mathrm{GW}}$ reconstruction from detector data..
@@ 4.1.1 Broken power law
We start considering a broken power law (BPL) template shown in the left panel of fig. 1 (see e.g. [43, 44, 45]):
@@
Our results are shown in fig. 2. We find that the scalar power spectrum $\mathcal{P}_{\zeta}$ can be accurately reconstructed near the peak, while larger uncertainties appear at the edges, reflecting the limitations of the input data. In particular, the constraints on the power spectrum in the infrared (IR) are almost entirely prior-dominated. This is due to two key reasons: i) IR modes far from the peak contribute negligibly to the SIGW spectrum around the peak, and ii) the IR tail of the SIGW spectrum generated during radiation domination exhibits a universal $k^{3}$ behaviour, up to logarithmic corrections (see e.g. [46, 47, 48, 12, 44, 49]). Nevertheless, it is known that in single field inflation $\mathcal{P}_{\zeta}$ can not normally growth steeper than $k^{4}$ in the infrared [50, 51] (but see e.g. [52, 53, 54, 55] for exceptions).
@@
These results can be understood analytically by noting that if $n_{\mathrm{IR}}$ and $n_{\mathrm{UV}}$ denote the IR and UV tilts of the curvature power spectrum, then the induced $\Omega_{\mathrm{GW}}$ spectrum exhibits the following asymptotic behavior (see e.g. [12]):
@@ 4.1.2 A template with oscillatory features
As a second representative template for the curvature power spectrum, we modify the BPL profile (4.3) including oscillatory features (middle column, fig. 1. See e.g. [56, 57, 58, 23] for physical motivations)
@@
The new parameters $A,\,B,\,C$ control the oscillations. We fix their values $A=10^{-3},B=15,C=2.5$ while the other parameters of eq. 4.6 have the same values as their BPL counterparts in eq. 4.4. The corresponding results are represented in fig. 4. The plot indicates that our Bayesian method is able to recover the shape of the oscillations quite well. Once again, the constraints in the tails are much weaker, especially in the IR since the IR part of the $\mathcal{P}_{\zeta}$ spectrum contributes very little to the SIGW spectrum for peaked $\mathcal{P}_{\zeta}$ spectra.
@@ 4.1.3 A sharply peaked spectrum
We next try to reconstruct a curvature spectrum characterized by a sharp feature, described by a modified log-normal profile  [^9] Log-normal scalar profiles have been much explored in the recent literature. See e.g. [23] for motivations and examples. (right column, fig. 1):
@@ 4.1.4 Non-detection
Finally, we present results for the case of a non-detection in fig. 6, where we only have upper limits on the $\Omega_{\mathrm{GW}}$ over a given frequency range. For a given detector, the upper limits on $\Omega_{\mathrm{GW}}$ would be the strongest around the peak sensitivity of the detector and weaker towards the ends, which is reflected in our mock upper limits shown in fig. 6. Since there is no detection in this case, we fix the number of nodes to four for simplicity although one can again obtain upper limits marginalised over models with different number of spline nodes as done in the previous examples. Our results demonstrate that this reconstruction method can also be applied to obtain upper limits on $\mathcal{P}_{\zeta}$ over a given frequency/wavenumber range which can complement the limits on $\mathcal{P}_{\zeta}$ obtained from PBH constraints. Given the exponential sensitivity of the PBH abundance to the perturbation amplitude [59], such limits can be very powerful in constraining the fraction of dark matter in PBH.
@@ 4.2 Early matter domination-radiation transition
In the previous section we focused on scenarios where primordial GWs are amplified during radiation domination upon re-entering the horizon. In this section and the next, we introduce an additional layer of complexity by considering cases where the universe is not necessarily radiation dominated at the time of GW formation. In particular, we consider a cosmological scenario in which inflation is followed by a period of matter domination that transitions quickly to RD, as laid out in section 2. Such an early matter domination (eMD) epoch is characterised by an equation of state parameter $w=0$, and we impose the relation $c_{\mathrm{s}}^{2}=w$ for the fluid sound speed [60].
@@
To calculate the induced GWs, we impose an abrupt cutoff on the power spectrum of curvature perturbations at $k_{\mathrm{cut}}$ which may be interpreted as the scale corresponding to the onset of eMD, or the scale where density perturbations become non-linear $k_{\mathrm{cut}}\sim 470/\eta_{\mathrm{rh}}$ [19]:
@@
Proceeding with additional tests, we perform for comparison a reconstruction assuming a pure radiation-dominated background using GW data generated under an early matter-dominated to radiation-dominated (eMD-RD) scenario for the BPL model. As shown in fig. 9, a resonance peak arising from the sudden reheating transition acts as a distinguishing feature between the two scenarios, with the Bayes factor decisively favouring the eMD-RD model. This indicates that, in principle, GW observations can be used to infer features of the transition. However, the enhancement in induced GWs is sensitive to the dynamics of the reheating process. While a sudden transition can lead to a pronounced enhancement, more gradual reheating, such as in models with constant decay rates [24] can suppress the signal, potentially weakening the apparent model preference. However, if the decay rate evolves in time, such that the field decays much more rapidly than the Hubble expansion rate at the time of transition, the induced GWs can still be significantly enhanced [61]. Thus, whether a bias in model reconstruction persists depends on the specific details of the transition dynamics, and further study is needed to assess the robustness of these conclusions across various reheating scenarios.
Additionally, it would be particularly interesting to explore whether the time dependence of the gravitational potential and its decay rate could be reconstructed from the GW signal itself. If this is feasible, it could provide insights into the underlying physics and thermal history, potentially allowing us to pinpoint key features such as (a) the onset of the eMD era, (b) its duration, and (c) the timescale of the transition to RD. Recent studies have begun to map the impact of these factors on $\Omega_{\text{GW}}$ [61, 24, 19, 62], and using the reconstruction methods proposed in this work, it might even be possible to directly infer this information from real signals, such as those detectable by LISA, offering a direct observational probe into the thermal history of the early universe, and possibly also the QCD transition (see e.g. [63, 64]). We leave these considerations for future study.
@@ 4.3 General equation of state
The transfer function $\mathcal{T}_{w}$ is given by eq. 2.18 [12]. Our aim in this case is to use GW data not only for reconstructing the momentum shape of the primordial power spectrum, but also the equation of state $w$ (we fix the sound speed such that $w=c_{s}^{2}$, i.e. treating this component as a perfect fluid). We make use of a modified version of the public code SIGWFast [30] to obtain the $\Omega_{\mathrm{GW}}(f)$ spectrum given an input curvature power spectrum, equation of state $w$ and transition/reheating scale $f_{\rm rh}$. SIGWFast calculates $\Omega_{\mathrm{GW}}(f)$ under the assumptions stated above (we take $f_{\rm rh}=10^{-5}\rm\,Hz$). [^10] Note that the code can only compute $\Omega_{\mathrm{GW}}(f)$ for $f>f_{\rm rh}$. We do not sample $f_{\rm rh}$ since it appears as an overall scaling factor in $\Omega_{\mathrm{GW}}(f)$, thus it is completely degenerate with the power spectrum amplitude. We leave an analysis of how $f_{\rm rh}$ could also be inferred by breaking this degeneracy, e.g. by additional theory motivated priors for $f_{\rm rh}$ or by requiring that the power spectrum does not overproduce PBH, to future studies.
@@
where $x$ and $x^{\prime}$ denote the positions of the nodes [^12] Not to be confused with $x=k\eta$ in section section 2., once again in $\log_{10}k$-space and $l$ corresponds to the GP lengthscale, which controls the correlation length of the GP predictions. Given positions of nodes $X$ and amplitudes $Y$, the prediction of the GP interpolation at a point $x_{*}$ is [65]
@@
Following this procedure, we obtain that both the primordial power spectrum as well as the equation of state are recovered fairly well – see fig. 11. However, we have larger uncertainties in the inferred $w$ as opposed to the sharply peaked case of fig. 10. This is due to the fact that due to the broad shape of the peak in $\mathcal{P}_{\zeta}$, the resonant peaks in $\Omega_{\mathrm{GW}}$ are absent. The previous discussion of section 4.1.1 with respect to the behaviour of the IR and UV tails also applies here, with only slight differences: eq. 4.5 now changes to [12]

### v3

@@ 3.2 Bayesian Model Selection via Bayesian Evidence
$\mathcal{Z}(\mathcal{M})=\int\mathcal{L}(d|\theta,\mathcal{M})\pi(\theta|\mathcal{M})\,d\theta,$ (3.1)
where $\mathcal{L}(d|\theta,\mathcal{M})$ is the likelihood of the data $d$ given the parameters $\theta$ and model $\mathcal{M}$, and $\pi(\theta|\mathcal{M})$ is the prior distribution of the model parameters [Trotta:2008qt]. Thus, for each spline model, the parameters $\theta$ over which we integrate are the spline node locations $\log_{10}k_{i}$, and the corresponding amplitudes $\log_{10}A_{i}$ of the scalar spectrum.
The Bayesian evidence automatically incorporates a penalty for excessive model complexity, effectively implementing Occam’s razor. In fact, a model with too many spline nodes has a larger parameter space to integrate over, diluting the integral and leading to a lower evidence unless the additional complexity is justified by the data. By comparing $\mathcal{Z}$ across different values of $N$, we select the models that best balance flexibility and predictive power. Typically, differences of $\Delta\log\mathcal{Z}\gtrsim 5$ between two models $(1,2)$ suggests that model $(1)$ is decisively favoured over model $(2)$, with $\Delta\log\mathcal{Z}_{(12)}=\log\mathcal{Z}_{(1)}-\log\mathcal{Z}_{(2)}$, while smaller differences ($\Delta\log\mathcal{Z}\lesssim 1$) are considered to be inconclusive in terms of Jeffrey’s scale [Trotta:2008qt].
The posterior distribution for the spline parameters can be highly complex, featuring multiple local maxima, as well as strong parameter degeneracies. Traditional sampling methods such as Markov Chain Monte Carlo (MCMC) struggle in such cases due to the difficulty of efficiently exploring the high-dimensional space. Additionally, local optimization methods are inadequate because they can become trapped in local extrema. For the same reasons, we opt not to use Akaike information criterion or other frequentist criteria for model comparison. To overcome these challenges, we employ nested sampling [SkillingNS], a particularly well-suited approach for computing Bayesian evidence while also generating posterior samples as a byproduct.
@@
$X(\bar{\mathcal{L}})=\int_{\mathcal{L}(\theta)>\bar{\mathcal{L}}}\pi(\theta)\,d\theta\,.$ (3.3)
Hence, $\bar{\mathcal{L}}$ can be thought of as the enclosing likelihood of cumulative prior mass $X$ which encompasses the region where the likelihood is larger than a threshold $\bar{\mathcal{L}}$. The core algorithm begins with a set of points drawn from the prior, called live points. At each step of the algorithm, the live point with the lowest likelihood is replaced by sampling from the prior, subject to the constraint that the likelihood at the new point must be larger than the current lowest likelihood value, and a statistical estimate of the prior mass $X$ is calculated. At the end of this process, we are left with a set of samples $\{\mathcal{L},X\}$ which can be used to estimate the evidence as well as approximate the posterior distribution of the parameters $\theta$ [2009BMC, Ashton:2022grj]. This method has several advantages over MCMC: in addition to providing an estimate of the evidence, it can efficiently handle multi-modal posteriors, and a complex posterior geometry. We implement nested sampling using the publicly available libraries PolyChord [Handley_2015, Handley:2015fda] and Nautilus [Lange:2023ydq].
@@ 3.3 Functional posteriors
$\mathbb{E}[\mathcal{P}_{\zeta,k}|\mathcal{M}]=\int d\theta\,\mathcal{P}_{\zeta,k}(\theta,\mathcal{M})p(\theta|d,\mathcal{M})\,$, (3.4)
@@ 4 Results
As an example, we choose the pivot frequency $f_{*}=10^{-3}\,\mathrm{Hz}$ as representative for LISA, with low uncertainties around the peak sensitivity $f_{*}$ and increasing ones towards the two ends of LISA’s sensitivity range (see e.g. [LISA:2024hlh]). We then use the result to create a Gaussian likelihood for the $\Omega_{\rm GW}$ as an input for the nested sampler, i.e.
@@
In our analysis, we assume that such $\Omega_{\mathrm{GW}}$ spectrum is reconstructed from detector data – e.g. using codes such as SGWBinner [Caprini:2019pxz, Flauger:2020qyi] in the case of LISA (see also [Hartwig:2023pft, Knapp:2025upw, Santini:2025iuj]) – thus providing datasets consisting of the observed values of $\Omega_{\mathrm{GW}}$ and their associated uncertainties $\Delta\Omega_{\mathrm{GW}}$, over certain frequency ranges specific to the instruments. [^6] This assumption helps us to focus our attention entirely on our reconstruction method but is not a necessity. This reconstruction method (or any other) can be easily incorporated into existing codes in such a manner that eliminates the need for a separate initial inference step to obtain $\Omega_{\mathrm{GW}},\,\Delta\Omega_{\mathrm{GW}}$. However, the two-step procedure has the advantage that the reconstruction method need not be applied repeatedly to the full dataset, but only to the $\Omega_{\rm GW}$ data, making it computationally more feasible.
While the spectra and error bars are highly idealised, and artificially generated with LISA as representative experiment to keep in mind, the methods we present do not rely on this being the case. They are applicable for any general form for $\Omega_{\mathrm{GW}}$ and any GW experiment. In particular, one may consider more realistic realizations of the $\Omega_{\mathrm{GW}}$ data simulated using the LISA noise model, without changing our method. [^7] In appendix C we also present results for noisy realisations of the $\Omega_{\mathrm{GW}}$ data. We should note though that the curvature spectrum reconstruction can only be as good as the $\Omega_{\mathrm{GW}}$ reconstruction from detector data.
@@ 4.1.1 Broken power law
We start considering a broken power law (BPL) template shown in the left panel of fig. 1 (see e.g. [Caprini:2024hue, LISACosmologyWorkingGroup:2024hsc, Marriott-Best:2024anh]):
@@
Our results are shown in fig. 2. The Bayesian evidence here favours the $N=4,5$ model the most and decreases as we increase the number of nodes, suggesting that the shape of the spectrum can be captured sufficiently well with 4 and 5 nodes. [^8] Nested sampling produces a statistical estimate of $\log\mathcal{Z}$, thus these numbers in fig. 2 come with their associated uncertainties. When running Nautilus to sample the posterior and compute the evidence, we choose settings to ensure that the resulting uncertainty is small $\Delta\log\mathcal{Z}<0.1$ (see [Lange:2023ydq] and also Nautilus FAQs) Since this is much smaller compared to the range over which $\log\mathcal{Z}$ varies as a function of the number of nodes, we do not show it in the plots.
We find that the scalar power spectrum $\mathcal{P}_{\zeta}$ can be accurately reconstructed near the peak, while larger uncertainties appear at the edges, reflecting the limitations of the input data. In particular, the constraints on the power spectrum in the infrared (IR) are almost entirely prior-dominated. This is due to two key reasons: i) IR modes far from the peak contribute negligibly to the SIGW spectrum around the peak, and ii) the IR tail of the SIGW spectrum generated during radiation domination exhibits a universal $k^{3}$ behaviour, up to logarithmic corrections (see e.g. [Caprini:2009fx, Cai:2019cdl, Yuan:2019wwo, Domenech:2021ztg, LISACosmologyWorkingGroup:2024hsc, Domenech:2025bvr]). Nevertheless, it is known that in single field inflation $\mathcal{P}_{\zeta}$ can not normally growth steeper than $k^{4}$ in the infrared [Byrnes:2018txb, Ozsoy:2019lyy] (but see e.g. [Carrilho:2019oqg, Tasinato:2020vdk, Tasinato:2023ukp, Cielo:2024poz] for exceptions).
@@
These results can be understood analytically by noting that if $n_{\mathrm{IR}}$ and $n_{\mathrm{UV}}$ denote the IR and UV tilts of the curvature power spectrum, then the induced $\Omega_{\mathrm{GW}}$ spectrum exhibits the following asymptotic behavior (see e.g. [Domenech:2021ztg]):
@@ 4.1.2 A template with oscillatory features
As a second representative template for the curvature power spectrum, we modify the BPL profile (4.3) including oscillatory features (middle column, fig. 1. See e.g. [Fumagalli:2021cel, Witkowski:2021raz, Braglia:2020taf, LISACosmologyWorkingGroup:2025vdz] for physical motivations)
@@
The new parameters A,B,CA,\,B,\,C control the oscillations. We fix their values $A=10^{-3},B=15,C=2.5$ while the other parameters of eq. 4.6 have the same values as their BPL counterparts in eq. 4.4. The corresponding results are represented in fig. 4. The plot indicates that our Bayesian method is able to recover the shape of the oscillations quite well. Once again, the constraints in the tails are much weaker, especially in the IR since the IR part of the $\mathcal{P}_{\zeta}$ spectrum contributes very little to the SIGW spectrum for peaked $\mathcal{P}_{\zeta}$ spectra.
@@ 4.1.3 A sharply peaked spectrum
We next try to reconstruct a curvature spectrum characterized by a sharp feature, described by a modified log-normal profile  [^10] Log-normal scalar profiles have been much explored in the recent literature. See e.g. [LISACosmologyWorkingGroup:2025vdz] for motivations and examples. (right column, fig. 1):
@@ 4.1.4 Non-detection
Finally, we present results for the case of a non-detection in fig. 6, where we only have upper limits on the $\Omega_{\mathrm{GW}}$ over a given frequency range. For a given detector, the upper limits on $\Omega_{\mathrm{GW}}$ would be the strongest around the peak sensitivity of the detector and weaker towards the ends, which is reflected in our mock upper limits shown in fig. 6. Since there is no detection in this case, we fix the number of nodes to four for simplicity although one can again obtain upper limits marginalised over models with different number of spline nodes as done in the previous examples. Our results demonstrate that this reconstruction method can also be applied to obtain upper limits on $\mathcal{P}_{\zeta}$ over a given frequency/wavenumber range which can complement the limits on $\mathcal{P}_{\zeta}$ obtained from PBH constraints. Given the exponential sensitivity of the PBH abundance to the perturbation amplitude [LISACosmologyWorkingGroup:2023njw], such limits can be very powerful in constraining the fraction of dark matter in PBH.
@@ 4.2 Early matter domination-radiation transition
In the previous section we focused on scenarios where primordial GWs are amplified during radiation domination upon re-entering the horizon. In this section and the next, we introduce an additional layer of complexity by considering cases where the universe is not necessarily radiation dominated at the time of GW formation. In particular, we consider a cosmological scenario in which inflation is followed by a period of matter domination that transitions quickly to RD, as laid out in section 2. Such an early matter domination (eMD) epoch is characterised by an equation of state parameter $w=0$, and we impose the relation $c_{\mathrm{s}}^{2}=w$ for the fluid sound speed [Assadullahi:2009nf].
@@
To calculate the induced GWs, we impose an abrupt cutoff on the power spectrum of curvature perturbations at $k_{\mathrm{cut}}$ which may be interpreted as the scale corresponding to the onset of eMD, or the scale where density perturbations become non-linear $k_{\mathrm{cut}}\sim 470/\eta_{\mathrm{rh}}$ [inomata_enhancement_2019]:
@@
Proceeding with additional tests, we perform for comparison a reconstruction assuming a pure radiation-dominated background using GW data generated under an early matter-dominated to radiation-dominated (eMD-RD) scenario for the BPL model. As shown in fig. 9, a resonance peak arising from the sudden reheating transition acts as a distinguishing feature between the two scenarios, with the Bayes factor decisively favouring the eMD-RD model. This indicates that, in principle, GW observations can be used to infer features of the transition. However, the enhancement in induced GWs is sensitive to the dynamics of the reheating process. While a sudden transition can lead to a pronounced enhancement, more gradual reheating, such as in models with constant decay rates [Inomata_gradual_2019] can suppress the signal, potentially weakening the apparent model preference. However, if the decay rate evolves in time, such that the field decays much more rapidly than the Hubble expansion rate at the time of transition, the induced GWs can still be significantly enhanced [pearce2024gravitationalwavesignalsearly]. Thus, whether a bias in model reconstruction persists depends on the specific details of the transition dynamics, and further study is needed to assess the robustness of these conclusions across various reheating scenarios.
Additionally, it would be particularly interesting to explore whether the time dependence of the gravitational potential and its decay rate could be reconstructed from the GW signal itself. If this is feasible, it could provide insights into the underlying physics and thermal history, potentially allowing us to pinpoint key features such as (a) the onset of the eMD era, (b) its duration, and (c) the timescale of the transition to RD. Recent studies have begun to map the impact of these factors on $\Omega_{\text{GW}}$ [pearce2024gravitationalwavesignalsearly, Inomata_gradual_2019, inomata_enhancement_2019, pearce2025usinggravitationalwavesignals], and using the reconstruction methods proposed in this work, it might even be possible to directly infer this information from real signals, such as those detectable by LISA, offering a direct observational probe into the thermal history of the early universe, and possibly also the QCD transition (see e.g. [Byrnes:2018clq, Franciolini:2023wjm]). We leave these considerations for future study.
@@ 4.3 General equation of state
The transfer function $\mathcal{T}_{w}$ is given by eq. 2.18 [Domenech:2021ztg]. Our aim in this case is to use GW data not only for reconstructing the momentum shape of the primordial power spectrum, but also the equation of state $w$ (we fix the sound speed such that $w=c_{s}^{2}$, i.e. treating this component as a perfect fluid). We make use of a modified version of the public code SIGWFast [Witkowski:2022mtg] to obtain the $\Omega_{\mathrm{GW}}(f)$ spectrum given an input curvature power spectrum, equation of state $w$ and transition/reheating scale $f_{\rm rh}$. SIGWFast calculates $\Omega_{\mathrm{GW}}(f)$ under the assumptions stated above (we take $f_{\rm rh}=10^{-5}\rm\,Hz$). [^11] Note that the code can only compute $\Omega_{\mathrm{GW}}(f)$ for $f>f_{\rm rh}$. We do not sample $f_{\rm rh}$ since it appears as an overall scaling factor in $\Omega_{\mathrm{GW}}(f)$, thus it is completely degenerate with the power spectrum amplitude. We leave an analysis of how $f_{\rm rh}$ could also be inferred by breaking this degeneracy, e.g. by additional theory motivated priors for $f_{\rm rh}$ or by requiring that the power spectrum does not overproduce PBH, to future studies.
@@
where $x$ and $x^{\prime}$ denote the positions of the nodes [^13] Not to be confused with $x=k\eta$ in section section 2., once again in $\log_{10}k$-space and $l$ corresponds to the GP lengthscale, which controls the correlation length of the GP predictions. Given positions of nodes $X$ and amplitudes $Y$, the prediction of the GP interpolation at a point $x_{*}$ is [williams2006gaussian]
@@
Following this procedure, we obtain that both the primordial power spectrum as well as the equation of state are recovered fairly well – see fig. 11. However, we have larger uncertainties in the inferred $w$ as opposed to the sharply peaked case of fig. 10. This is due to the fact that due to the broad shape of the peak in $\mathcal{P}_{\zeta}$, the resonant peaks in $\Omega_{\mathrm{GW}}$ are absent. The previous discussion of section 4.1.1 with respect to the behaviour of the IR and UV tails also applies here, with only slight differences: eq. 4.5 now changes to [Domenech:2021ztg]